Use the quadratic formula to solve the following quadratic equations, giving your answer to 2 decimal places in each case.
X2+6x+4=0
3x2- 6x+2=0
6x2-9x-4=0
The 2 after the x is squared

Answer by TimothyLamb(4379)

xx + 6x + 4 = 0
3xx - 6x + 2 = 0
6xx - 9x - 4 = 0
---
All of the above quadratic equations are in standard form.
I will solve the first one, you do the rest using the same technique:
---
the first equation has a=1, b=6, and c=4
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 6 4
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = -0.763932023
x = -5.23606798
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php